# Process-based Indicators of Vulnerability Re-Introducing Code Changes: An Exploratory Case Study

### v1

@@ Abstract.
Software vulnerabilities often persist or re-emerge even after being fixed, revealing the complex interplay between code evolution and socio-technical factors. While source code metrics provide useful indicators of vulnerabilities, software engineering process metrics can uncover patterns that lead to their introduction. Yet few studies have explored whether process metrics can reveal risky development activities over time — insights that are essential for anticipating and mitigating software vulnerabilities. This work highlights the critical role of process metrics along with code changes in understanding and mitigating vulnerability reintroduction. We move beyond file-level prediction and instead analyze security fixes at the commit level, focusing not only on whether a single fix introduces a vulnerability but also on the longer sequences of changes through which vulnerabilities evolve and re-emerge. Our approach emphasizes that reintroduction is rarely the result of one isolated action, but emerges from cumulative development activities and socio-technical conditions. To support this analysis, we conducted a case study on the ImageMagick project by correlating longitudinal process metrics such as bus factor, issue density, and issue spoilage with vulnerability reintroduction activities, encompassing 76 instances of reintroduced vulnerabilities. Our findings show that reintroductions often align with increased issue spoilage and fluctuating issue density, reflecting short-term inefficiencies in issue management and team responsiveness. These observations provide a foundation for broader studies that combine process and code metrics to predict risky fixes and strengthen software security.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software testing and debuggingccs: Security and privacy Software security engineering
@@ 1. Introduction
Figure 1. Overview of the vulnerability reintroduction process. Contributors submit code to an open-source project which undergoes a software review process before being committed to the version control system. After the commit is made, a vulnerability in the contribution may be identified and disclosed to a vulnerability database (e.g., CVE, NVD). After which contributors submit new code that addresses the vulnerability while simultaneously reintroducing a vulnerability. This reintroduced vulnerability is then identified where it is patched out by a contributor in a potentially cyclical pattern.
@@
Software security remains a critical challenge within open-source software. The rapid pace and complexity of technological advancement places software engineers under time constraints, increasing the risk of flawed implementations and introducing vulnerabilities. Software vulnerabilities are reported to vulnerability databases including the Common Vulnerabilities and Exposures (CVE) (The MITRE Corporation, 2025b) and National Vulnerability Database (NVD) (National Institute Of Standards And Technology, 2025).
@@
Despite software engineers’ efforts to address vulnerabilities, their fixes can inadvertently introduce new security issues (Śliwerski et al., 2005; Shimmi and Rahimi, 2022) as visualized in Figure˜1. Vulnerable software systems compromise overall system dependability (Braz et al., 2022), hampers reproducibility (Kaur et al., 2021), and diminishes trust (Wermke et al., 2022) Moreover, the integration of software dependencies packages enables vulnerabilities to propagate through a software supply chain (Zimmermann et al., 2019; Zahan et al., 2022; Hegewald and Beyer, 2025). Although automated security audits (Zahan et al., 2023; Snyk Limited, 2025; Google, 2025; GitHub, 2025a) and tooling (Nong et al., 2023; Alazmi and De Leon, 2022; Nie et al., 2022; Sheng et al., 2025) are available to the open-source community, the impact of these tools on engineering decisions is limited (Eisty et al., 2018; Ayala et al., 2025). While research has examined software vulnerability prediction (Nong et al., 2023; Sheng et al., 2025), a gap persists in identifying vulnerability reintroduction activities where remediation of one vulnerability inadvertently introduces another (Shimmi and Rahimi, 2022). Furthermore, as vulnerabilities are introduced and reintroduced as commit operations, a gap exists in analyzing the software engineering process involved in such operations.
@@
To address this, we conduct the first empirical case study  of a software’s engineering process and its correlation to software vulnerability reintroduction activities. We identify pairs of vulnerability reintroducing and fixing commits and automatically compute longitudinal process metrics for bus factor, issue density, and issue spoilage for the ImageMagick (ImageMagick Studio LLC, 2024) project — a well known image processing project. Our results show that (1) ImageMagick has maintained a healthy number of contributors, implying that there is always a contributor available to fix and review vulnerability reintroducing commits; (2) ImageMagick’s issue density has remained low even during periods of vulnerability reintroduction implying that maintainers can adequately handle incoming issues while working on new features; implying that the project’s development workflow can accommodate backlog accumulation and ensuring that regular development activities are not disrupted by recurring vulnerabilities. (3) ImageMagicks’s issue spoilage typically rises during the time window of reintroduced vulnerabilities, implying that contributors are not closing issues as fast as they are opened. Periods characterized by vulnerability reintroductions coincide with increased issue spoilage, suggesting that developers may temporarily shift focus toward remediation efforts at the expense of issue closure rates. This pattern highlights a potential process bottleneck during security-intensive phases, where parallel task management may strain available resources.
@@
Taken together, these findings suggest that while ImageMagick’s contributor base and issue management processes provide a strong foundation for sustainable maintenance, vulnerability reintroductions still impose measurable strain on project dynamics. The increase in issue spoilage during such periods indicates that even well-staffed open-source projects face coordination and prioritization challenges when balancing security remediation with ongoing development. This underscores the need for improved process-aware tooling and decision support systems that help developers anticipate, detect, and mitigate reintroduction risks before they propagate through the codebase.
Our contributions are:
  1. (1)
@@
    The first empirical case study correlating longitudinal software engineering process metrics with open-source software engineering vulnerability reintroduction activities.
  2. (2)
    A methodology identify for identifying vulnerability reintroducing commits leveraging existing algorithms (Śliwerski et al., 2005) and LLMs.
  3. (3)
    A dataset of 76 vulnerability-reintroducing commits that extends existing vulnerability datasets by including future commits that fix the reintroduced vulnerabilities.
@@ 2. Background and Related Work
In this section, we give a brief overview software engineering process metrics, process metrics for predicting software vulnerabilities, and software vulnerability reintroduction activities.
@@ 2.1. Software Engineering Process Metrics
Software engineering process metrics quantify the activity involved in the creation or maintenance of a software system (Eisty et al., 2018). Prior work has shown that code complexity, churn, and developer activity metrics can serve as indicators of software vulnerabilities (Shin et al., 2010) however the open-source community has seen limited adoption of these and other process metrics (Eisty et al., 2018). However, these approaches primarily rely on measurements taken at specific releases, rather than examining longitudinal trends. Captured at a specific release, snapshot process metrics offer insights into the software system’s current state, however the measuring process metrics longitudinally captures socio-technical trends influencing project evolution (Synovic et al., 2022). Several process metrics have been proposed (Fritz et al., 2014; Goggins et al., 2021; Linux Foundation, [n. d.]), but for our study we focus on derived process metrics including bus factor, issue density, and issue spoilage.
Bus Factor measures team-wide familiarity with a system’s components by counting contributions to various elements (Cosentino et al., 2015); high bus factors indicate widespread understanding, while a low bus factor (i.e., less than one (Piggot and Amrit, 2013)) suggests siloed knowledge potentially limiting the number of active contributors who could resolve vulnerabilities.
Issue Density extends the defect density metric (Fenton and Bieman, 2014) to open-source issue trackers; high density suggests engineers struggle to manage community submissions efficiently, whereas low density implies timely responses to these inputs.
Issue spoilage quantifies how quickly engineering teams close reported issues; high spoilage indicates unaddressed lingering issues, whereas low spoilage signifies prompt handling of issues.
Our study advances software engineering process metrics by analyzing longitudinal process metrics, including bus factor, issue density, and issue spoilage, to correlate engineering activities vulnerability reintroduction. This approach underscores the critical gap between awareness and practical adoption of such metrics in software engineering communities, emphasizing their potential as actionable indicators for improving code quality and security resilience.
@@ 2.2. Software Vulnerabilities and Vulnerability Reintroduction
A software vulnerability refers to specific weakness in a software system’s code that can be exploited to compromise the confidentiality, integrity, or availability of a system. Centralized databases aggregate, report, score, and distribute publicly disclosed vulnerabilities to enable vulnerability tracking and awareness (National Institute Of Standards And Technology, 2025; The MITRE Corporation, 2025b; Google, 2025; GitHub, 2025b; GitLab, 2025). Vulnerabilities can be classified via the Common Weakness Enumeration (CWE) schema (The MITRE Corporation, 2025b) which captures recurring weaknesses in software (e.g., buffer overflows, improper input validation) that automated tooling can classify (Das et al., 2021; Han et al., 2017). Furthermore the Common Vulnerability Scoring System (CVSS) provides a means to quantify the severity of a vulnerability from a range of zero to ten (Forum of Incident Response and Security Teams, 2024), with ten being reserved for the most severe vulnerabilities (Benny Isaacs et al., 2025; Chen Zhaojun, 2021).
Software fixes — applied commits — attempt to address and remove the software vulnerability, but previous work has found that not all fixes are permanent (Śliwerski et al., 2005; Braz et al., 2022). Prior work has identified that vulnerability reintroduction occurs when a fix inadvertently reintroduces a new security weakness as confirmed by later fixes (Shimmi and Rahimi, 2022; Braz et al., 2022). Software vulnerability reintroduction was conceptually identified as a threat to software systems via attack–defense co-evolution by positing that a vulnerability fixing operation might itself serve as a predictor of a subsequent vulnerability (Shimmi and Rahimi, 2022). Such reintroductions pose significant risks because they may silently undo prior security improvements and leave systems exposed once again.
While existing datasets have labeled vulnerable and non-vulnerable source code (Chen et al., 2023; Fan et al., 2020), to our knowledge, there are no datasets that specifically capture vulnerability reintroducing commits and their fixes. Our work resolves this by releasing a dataset that extends established vulnerability datasets (Chen et al., 2023; Fan et al., 2020) with 81 pairs for the ImageMagick project that both resolve existing CWEs while simultaneously reintroducing new vulnerabilities.
@@ 3. Case Study: ImageMagick
In this section we introduce our case study on ImageMagick — a well known and embeddable image processing project. We discuss an overview of the problem, our methodology to evaluate the project’s software engineering process metrics and high CVSS scoring CVEs longitudinally, and analyze the data collected during our case study. We selected ImageMagick as our case study due to it being a well known open-source image processing project with a significant development history (ImageMagick Studio LLC, 2024), written primarily in C potentially enabling common weaknesses (e.g., buffer overflows, memory leaks), and being readily available in existing vulnerability datasets that have previously mapped CVE disclosures to specific commits (Chen et al., 2023; Fan et al., 2020).
@@ 3.1. Problem Overview
In this section we show an example of a vulnerability that was reintroduced in the fix for a separate vulnerability. CVE-2018-11625 (The MITRE Corporation, 2018a) was an out of bounds read vulnerability (The MITRE Corporation, 2025c) with a CVSS severity score of 8.8 disclosed on May 31st, 2018.
Vulnerability Reintroducing Fix
Listing 1 shows a commit accepted by ImageMagick that reintroduced a vulnerability while attempting to resolve this vulnerability. On May 30th, 2018 commit 5294966 was accepted into ImageMagick to resolve a buffer overflow by allocating an additional byte for colormap_index. While the intention was to prevent out-of-bounds writes, this fix was potentially unsafe when image->colors or MaxColormapSize was smaller than MaxMap, creating scenarios where the buffer could still overflow under certain image configurations. [^1] https://github.com/ImageMagick/ImageMagick/commit/5294966
@@
Reintroduced Vulnerability Fix
Listing 2 shows the subsequent commit c111ed9 that corrected the reintroduced vulnerability on April 8th, 2019. This issue was resolved by ensuring that the allocation size was properly bounded using the MagickMax() function, guaranteeing that the memory allocated for colormap_index was at least MaxMap. This change prevented buffer overflow conditions that could occur when the color count was smaller than the maximum mapping size. [^2] https://github.com/ImageMagick/ImageMagick/commit/280215b9936d145dd5ee91403738ccce1333cab1
The reintroduced vulnerability remained active for 313 days across 1,094 commits and 6 releases.
@@
Figure 2. Our methodology to extract and compute longitudinal SEP metrics for vulnerability reintroducing and fixing commits in the ImageMagick project. We extract issues and pull requests from the GitHub issue tracker for ImageMagick to compute issue density and issue spoilage. Git commits are leveraged to compute bus factor and issue density, and to identify vulnerability fixing commits with the SZZ algorithm and Anthropic Claude 3.5 Sonnet from the ImageMagick GitHub repository, BigVul, and DiverseVul. Data is also collected from CVE and is leveraged to visualize the data.
@@ 3.2. Methodology
To evaluate the socio-technical context in which vulnerabilities are reintroduced into the ImageMagick dataset we extract and evaluate vulnerability resolving commits from existing datasets, extend existing algorithms and methods with LLMs to identify pairs of reintroducing and fixing commits, and leverage prior work on evaluating longitudinal software engineering process metrics on selected high CVSS scoring CVEs that were identified to reintroduce issues. Our goal is twofold: (1) to flag vulnerability fixing commits that are likely to trigger regressions, and (2) to isolate the engineering process patterns that contribute to vulnerability reintroducing activity in open-source software. Our methodology is outlined in Figure˜2

### v2

@@ Abstract.
Software vulnerabilities often persist or re-emerge after being fixed, revealing the complex interplay between code evolution and socio-technical factors. While source code metrics provide useful indicators of vulnerabilities, software engineering process metrics can measure engineering activities that lead to their introduction. To our knowledge, few studies have explored whether these metrics can identify these activities over time — insights that are essential for anticipating and mitigating software vulnerabilities. This work highlights the critical role of process metrics along with code changes in understanding and mitigating vulnerability reintroduction. We move beyond file-level prediction and instead analyze vulnerability fixes at the commit level, focusing on sequences of changes through which vulnerabilities re-emerge.
Our approach emphasizes that reintroduction is rarely the result of one isolated action, but emerges from cumulative engineering activities and socio-technical conditions. To support this analysis, we conduct a case study on the ImageMagick project by correlating longitudinal process metrics including bus factor, issue density, and issue spoilage with vulnerability reintroduction activities, encompassing 76 instances of reintroduced vulnerabilities. Our findings show that vulnerability reintroduction often aligns with increased issue spoilage and fluctuating issue density, reflecting short-term inefficiencies in issue management and team responsiveness. These observations provide a foundation for broader studies that combine process and code metrics to predict vulnerability reintroducing fixes and strengthen software security.
@@
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software testing and debuggingccs: Security and privacy Software security engineering
Figure 1. Overview of the vulnerability reintroduction process. Contributors submit code that passes review and enters the version control system. Once a vulnerability is disclosed and patched, a new flaw is inadvertently reintroduced in the fix, leading to a recurring cycle of identification and patching.
@@ 1. Introduction
Software security remains a critical challenge within open-source software. The rapid pace and complexity of technological advancement place software engineers under time constraints, increasing the risk of flawed implementations and introducing vulnerabilities. Software vulnerabilities are reported to the National Vulnerability Database (NVD) (National Institute Of Standards And Technology, 2025) and Common Vulnerabilities and Exposures (CVE) (The MITRE Corporation, 2025b) databases.
Despite software engineers’ efforts to address vulnerabilities, their fixes can inadvertently introduce new security issues (Śliwerski et al., 2005; Shimmi and Rahimi, 2022) as visualized in Figure˜1. Vulnerable software systems compromise overall system dependability (Braz et al., 2022), hamper reproducibility (Kaur et al., 2021), and diminish trust (Wermke et al., 2022)
@@
Moreover, the integration of software dependencies enables vulnerabilities to propagate through software supply chains (Zimmermann et al., 2019; Zahan et al., 2022; Hegewald and Beyer, 2025). Although automated security audits (Zahan et al., 2023; Snyk Limited, 2025; Google, 2025; GitHub, 2025a) and tooling (Nong et al., 2023; Alazmi and De Leon, 2022; Nie et al., 2022; Sheng et al., 2025) are available to the open-source community, the impact of these tools on engineering decisions is limited (Eisty et al., 2018; Ayala et al., 2025). While research has examined software vulnerability prediction (Nong et al., 2023; Sheng et al., 2025), to our knowledge, a gap persists in identifying vulnerability reintroducing activities where remediation of one vulnerability inadvertently introduces another (Shimmi and Rahimi, 2022). Furthermore, as vulnerabilities propagate through commit operations, a gap exists in analyzing the software engineering process involved.
@@
To address this, we conduct the first exploratory case study of a software’s engineering process and its correlation to vulnerability reintroduction. We identify pairs of vulnerability reintroducing and fixing commits and automatically compute longitudinal process metrics for bus factor, issue density, and issue spoilage for the ImageMagick (ImageMagick Studio LLC, 2024) project — a well known image processing project. Our results show that (1) ImageMagick has maintained a healthy number of contributors, implying that there is always a contributor available to fix and review vulnerability reintroducing commits; (2) ImageMagick’s issue density has stayed low even during vulnerability reintroductions, demonstrating that maintainers can manage incoming issues efficiently while developing new features, suggesting the project’s workflow can handle backlog accumulation without disrupting regular development activities; and (3) mageMagick’s issue backlog spikes during vulnerability reintroductions, suggesting that contributors prioritize remediation over routine issue closure. This shift in focus toward security fixes creates a bottleneck in parallel task management, straining resources during periods of high-security workloads.
@@
Taken together, these findings suggest that while ImageMagick’s contributor base and issue management processes provide a strong foundation for sustainable maintenance, vulnerability reintroductions still impose measurable strain on project dynamics. This underscores the need for improved process-aware tooling and decision support systems that help engineers anticipate, detect, and mitigate reintroduction risks before they propagate through the codebase.
@@
Our contributions are: (1)The first empirical case study correlating longitudinal software engineering process metrics with open-source software engineering vulnerability reintroduction activities; (2) A methodology for identifying vulnerability reintroducing commits leveraging existing algorithms (Śliwerski et al., 2005) and LLMs; and (3) A dataset of 76 vulnerability-reintroducing commits that extends existing vulnerability datasets by including future commits that fix the reintroduced vulnerabilities.
We release our source code and dataset artifacts on Zenodo (Shimmi et al., 2026).
@@ 2. Background and Related Work
In this section, we give a brief overview of software engineering process metrics, process metrics for predicting software vulnerabilities, and vulnerability reintroduction.
@@ 2.1. Software Engineering Process Metrics
Software engineering process metrics quantify the activity involved in the creation or maintenance of a software system (Eisty et al., 2018). Prior work has shown that code complexity, churn, and developer activity metrics can serve as indicators of software vulnerabilities (Shin et al., 2010), however the open-source community has seen limited adoption of these (Eisty et al., 2018). These metrics are computed at specific releases, rather than examining longitudinal trends. Captured at a specific release, snapshot process metrics offer insights into the software system’s current state, however, measuring process metrics longitudinally captures socio-technical trends influencing project evolution (Synovic et al., 2022). Several process metrics have been proposed (Fritz et al., 2014; Goggins et al., 2021; Linux Foundation, 2025), but for our study, we focus on derived process metrics including bus factor, issue density, and issue spoilage.
Bus Factor measures team-wide familiarity with a system’s components by counting contributions to various elements (Cosentino et al., 2015); high bus factors indicate widespread understanding, while a low bus factor (i.e., less than one (Piggot and Amrit, 2013)) suggests siloed knowledge potentially limiting the number of active contributors who could resolve vulnerabilities. Issue Density extends the defect density metric (Fenton and Bieman, 201) to open-source issue trackers; high density suggests engineers struggle to manage community submissions efficiently, whereas low density implies timely responses to these inputs. Issue spoilage quantifies how quickly engineering teams close reported issues; high spoilage indicates unaddressed lingering issues, whereas low spoilage signifies prompt handling of issues. A detailed overview of each of these metrics is presented in (Thiruvathukal et al., 2018; Synovic et al., 2022).
Our study advances software engineering process metrics by analyzing several process metrics longitudinally to correlate engineering activities to vulnerability reintroduction. This approach underscores the critical gap between awareness and practical adoption of such metrics in software engineering communities, emphasizing their potential as actionable indicators for improving code quality and security resilience.
@@ 2.2. Software Vulnerabilities and Vulnerability Reintroduction
A software vulnerability refers to specific weakness in a software system’s code that can be exploited to compromise the confidentiality, integrity, or availability of a system. Centralized databases aggregate, report, score, and distribute publicly disclosed vulnerabilities to enable vulnerability tracking and awareness (National Institute Of Standards And Technology, 2025; The MITRE Corporation, 2025b; Google, 2025; GitHub, 2025b; GitLab, 2025). Vulnerabilities can be classified via the Common Weakness Enumeration (CWE) schema (The MITRE Corporation, 2025b) which captures recurring weaknesses in software (e.g., buffer overflows, improper input validation) that automated tooling can classify (Das et al., 2021; Han et al., 2017). Furthermore, the Common Vulnerability Scoring System (CVSS) provides a means to quantify the severity of a vulnerability from a range of zero to ten (Forum of Incident Response and Security Teams, 2024), with ten being reserved for the most severe vulnerabilities (Benny Isaacs et al., 2025; Chen Zhaojun, 2021).
Software fixes attempt to address and remove software vulnerabilities, but prior work has found that not all fixes are permanent (Śliwerski et al., 2005; Braz et al., 2022). Vulnerability reintroduction occurs when a fix inadvertently reintroduces a new security weakness as confirmed by later fixes (Shimmi and Rahimi, 2022; Braz et al., 2022). Software vulnerability reintroduction was conceptually identified as a threat to software systems via attack–defense co-evolution by positing that a vulnerability fixing operation might itself serve as a predictor of a subsequent vulnerability (Shimmi and Rahimi, 2022). Such reintroductions pose significant risks because they may silently undo prior security improvements and leave systems exposed once again.
While existing datasets have labeled vulnerable and non-vulnerable source code (Chen et al., 2023; Fan et al., 2020), to our knowledge, there are no datasets that specifically capture sequences of commits that resolve and reintroduce vulnerabilities. Our work resolves this by releasing a dataset that extends established vulnerability datasets (Chen et al., 2023; Fan et al., 2020) with 81 pairs for the ImageMagick project that both resolve existing CWEs while simultaneously reintroducing new vulnerabilities.
@@ 3. Case Study: ImageMagick
In this section, we introduce our case study on ImageMagick — a well-known image processing project. We discuss an overview of the problem, our methodology to evaluate the project’s software engineering process metrics and high CVSS scoring CVEs longitudinally, and analyze the data collected during our case study. We selected ImageMagick as our case study due to it being a well-known open-source image processing project with a significant development history (ImageMagick Studio LLC, 2024), written primarily in C potentially enabling common weaknesses (e.g., buffer overflows, memory leaks), and being readily available in existing vulnerability datasets that have previously mapped CVE disclosures to specific commits (Chen et al., 2023; Fan et al., 2020).
@@ 3.1. Problem Overview
In this section we show an example of a vulnerability being reintroduced as a side effect for resolving a different vulnerability. This vulnerability remained active for 313 days across 1,094 commits and 6 releases.
Vulnerability Reintroducing Fix CVE-2018-11625 (The MITRE Corporation, 2018a) was an out of bounds read vulnerability (The MITRE Corporation, 2025c) with a CVSS severity score of 8.8 disclosed on May 31st, 2018. Listing 1 shows a commit accepted by ImageMagick that reintroduced a vulnerability while attempting to resolve CVE-2018-11625. On May 30th, 2018 commit 5294966 was accepted into ImageMagick to resolve a buffer overflow by allocating an additional byte for colormap_index. While the intention was to prevent out-of-bounds writes, this fix was potentially unsafe when image->colors or MaxColormapSize was smaller than MaxMap, creating scenarios where the buffer could still overflow under certain image configurations. [^1] https://github.com/ImageMagick/ImageMagick/commit/5294966
@@
Reintroduced Vulnerability Fix Listing 2 shows a subsequent commit c111ed9 that corrected the reintroduced vulnerability on April 8th, 2019. This issue was resolved by ensuring that the allocation size was properly bounded using the MagickMax() function, guaranteeing that the memory allocated for colormap_index was at least MaxMap. This change prevented buffer overflow conditions that could occur when the color count was smaller than the maximum mapping size. [^2] https://github.com/ImageMagick/ImageMagick/commit/280215b9936d145dd5ee91403738ccce1333cab1
@@
Figure 2. We analyzed the ImageMagick project by extracting issues and pull requests from GitHub to calculate issue density and spoilage. Using Git commits from GitHub, BigVul, and DiverseVul, we computed the bus factor and identified vulnerability-fixing commits via the SZZ algorithm and Claude 3.5 Sonnet. Finally, CVE data was integrated to enrich the analysis and facilitate data visualization.
@@ 3.2. Methodology
To evaluate the socio-technical context in which vulnerabilities are reintroduced into ImageMagick, we extract and evaluate vulnerability resolving commits from two existing datasets, extend existing algorithms and methods with LLMs to identify pairs of reintroducing and fixing commits, and leverage prior work on evaluating longitudinal software engineering process metrics on selected high CVSS scoring CVEs that were identified to reintroduce issues. Our goal is twofold: (1) to flag vulnerability fixing commits that are likely to trigger regressions, and (2) to isolate the engineering process patterns that contribute to vulnerability reintroducing activity in open-source software. Our methodology is outlined in Figure˜2